# AI-Powered Autonomous Weapons Risk Geopolitical Instability and Threaten AI Research

### v1

@@ 4.1 Historical AWS Policy Efforts
To begin this section, we will survey a few major events in AWS regulation and policy over the past two decades. Awareness and concern about AWS began to increase after 2001, when the US Congress ordered the Pentagon to make one-third of their “deep strike” aircraft fleet unmanned by 2010, and one-third of their ground combat vehicles unmanned by 2015 (U.S. Congress, 2001). These timelines proved overly aggressive, but it sparked a number of efforts to regulate AWS development: In 2010, the International Committee for Robot Arms Control called for regulations on remote-controlled and autonomous weapons (Markoff, 2010). In 2013, the Campaign to Stop Killer Robots started as a coalition of NGOs attempting to push for global bans on lethal AWS (Stop Killer Robots, 2021). In 2014, the phrase “meaningful human control” (MHC) was coined by Article 36, which has since become a core principle of AWS ethics and regulation (Article 36, 2014).
Calls for regulation have not been limited to NGOs—A number of national and international efforts to regulate AWS have been seen in recent years. At the national level, the Block Nuclear Launch by Autonomous Artificial Intelligence Act was introduced in 2023 in the US (Markey, 2023). Internationally, in 2023 60 countries, including China and the USA, issued a call to action in the Hague towards more responsible use of AI for military applications (Sterling, 2023). Also in 2023, the UN passed a resolution calling for strict regulation on AWS, and for humans to stay in control of lethal decision-making in war (UN, 2023). Lastly, while not directly targeting AWS, The EU has moved to regulate AI risks, first in 2021 and more recently in 2024, with the EU AI Act set to pass as the first major regulation on artificial intelligence, though its implications for AWS are currently unclear (European Parliament News, 2023).
@@ 4.2 International Policy Recommendations
The first policy recommendation we make is that nations can mitigate the negative effects of AWS on conflict risk by maintaining a ban on human-independent use of AWS. That is, AWS should not be deployed into combat without a meaningful human battlefield presence alongside them. While the widespread deployment of AWS is likely inevitable, there is still time to influence where and how it is deployed, and human-independent deployment is not technically trivial.
There are two avenues to preventing human-independent use: AWS manufacturers should not develop systems designed to operate in an environment free of friendly humans, and policymakers should not field independent combat formations composed entirely of AWS.On the technical front, requiring strong human-in-the-loop measures goes a long way, as does limiting the use of generalist AI models in AWS. The AI research community should be wary of and resist attempts to use large pretrained models in AWS development. These models are likely more useful for building generalist AWS capable of human-independent operations compared to more specialized AWS, which can be developed using smaller and more specialized ML models. Further, AWS using specialist models will likely prove easier to develop to maturity, and more easily fit into existing military command systems and doctrine.
@@ 4.3 Academic Policy and Action Recommendations
Lastly, we present several actions for academics and academic institutions to take to address the issues we describe in this work, beginning with an overview of recent actions academics have taken on AWS regulation. Most recent academic initiatives have centered on open letters or petitions seeking to regulate or ban AWS (Future of Life Institute, 2016; Amoroso & Tamburrini, 2020). For example, in 2017 a Clearwater Robotics co-founder published an open letter calling for the UN to ban AWS, with 116 founders of robotics and AI companies from 26 countries signing on (Bogdon, 2017). Similarly, in 2018 AI researchers from 30 countries wrote an open letter protesting a new department at KAIST that merged academic and military research (Walsh, 2018). In 2018, Google employees’ protest led to the company canceling a Pentagon contract that could have led to more accurate drone strike technology (Crofts, P. & van Rijswijk, H., 2020). Another Google employee protest has started recently, opposing their Project Nimbus (Perrigo, 2024). Furthermore, we note that global public opinion is favorable for AWS regulation currently—A 2020-2021 public opinion survey found that approximately 61 percent of adults across 28 countries oppose the use of fully autonomous weapons, with India being the only country surveyed with majority support for AWS (Ipsos, 2021).
@@ Acknowledgements
Last, some may argue that AWS will become essential for national security, and thus that it is unethical to discourage nations from developing AWS because they are needed to maintain military deterrence. More broadly, one could ask under what extremes may highly-autonomous AWS development and production be viewed as a necessary evil, and discouraging it therefore unethical? As a hypothetical example, an aggressive world power may refuse to limit fully-autonomous/human-free AWS development, and over years or decades their technology could progress to the point where it is impossible to counter using conventional forces, human-controlled AWS, or nuclear deterrence. In this particular situation we concede that fully-autonomous AWS development with a deterrence mentality, mirroring modern nuclear deterrence and tenuous geopolitical equilibriums, may pragmatically become a necessity. However, as outlined in this paper, while AWS battlefield deployments have started in preliminary ways, AWS are currently not the core of any nation’s military, and global public and institutional opinions are largely in favor of substantial regulation of some form. As such, this hypothetical is far from guaranteed to come to pass. However, we do not believe that the contributions of this paper are contingent on taking a position regarding the above hypothetical. Our work concerns the impacts of AWS on conflict risk and academic freedom, which will occur regardless of whether AWS are under strict human control or are highly autonomous, so long as they are critical to warfare and replace the human battlefield element. The goal of this work is to open discussion on the impacts of AWS among AI researchers and ground it in prior work, and we expect that the ethics of meaningful human control will be a major topic of future debate in the community, but our contribution does not rely on taking an ethical position on the matter.

### v2

@@ 4.1 Historical AWS Policy Efforts
To begin this section, we will survey a few major events in AWS regulation and policy over the past two decades. Awareness and concern about AWS began to increase after 2001, when the US Congress ordered the Pentagon to make one-third of their “deep strike” aircraft fleet unmanned by 2010, and one-third of their ground combat vehicles unmanned by 2015 (U.S. Congress, 2001). These timelines proved overly aggressive, but sparked a number of efforts to regulate AWS development—In 2010, the International Committee for Robot Arms Control called for regulations on remote-controlled and autonomous weapons (Markoff, 2010). Then, in 2013, the Campaign to Stop Killer Robots started as a coalition of NGOs attempting to push for global bans on lethal AWS (Stop Killer Robots, 2021). In 2014, the phrase “meaningful human control” (MHC) was coined by Article 36, which has since become a core principle of AWS ethics and regulation (Article 36, 2014).
Calls for regulation have not been limited to NGOs—A number of national and international efforts to regulate AWS have been seen in recent years. As of 2020, at least 30 countries have called for a ban on lethal AWS, with many others expressing some degree of concern about AWS (Stauffer, 2020). In the USA, the Block Nuclear Launch by Autonomous Artificial Intelligence Act was introduced in 2023 (Markey, 2023). In 2023 60 countries, including China and the USA, issued a call to action in the Hague towards more responsible use of AI for military applications (Sterling, 2023). Also in 2023, the UN passed a resolution calling for strict regulation on AWS, and for humans to stay in control of lethal decision-making in war (UN, 2023). Lastly, while not directly targeting AWS, the EU has moved to regulate AI risks, first in 2021 and more recently in 2024, with the EU AI Act passed as the first major regulation on artificial intelligence, though its implications for AWS are currently unclear (European Parliament News, 2023).
@@ 4.2 International Policy Recommendations
The first policy recommendation we make is that nations can mitigate the negative effects of AWS on conflict risk by maintaining a ban on human-independent use of AWS. That is, AWS should not be deployed into combat without a meaningful human battlefield presence alongside them, whether using human-AWS teaming processes or as part of separate operations in the same battlespace. While the widespread deployment of AWS is likely inevitable, there is still time to influence where and how it is deployed, and human-independent deployment is not technically trivial.
There are two avenues to preventing human-independent use: AWS manufacturers should not develop systems designed to operate in an environment free of friendly humans, and policymakers should not field independent combat formations composed entirely of AWS. On the technical front, requiring strong human-in-the-loop measures goes a long way, as does limiting the use of generalist AI models in AWS. The AI research community should be wary of and resist attempts to use large pretrained models in AWS development—These models are likely more useful for building generalist AWS capable of human-independent operations compared to more specialized AWS, which can be developed using smaller and more specialized ML models. Further, AWS using specialist models will likely prove easier to develop to maturity and more easily fit into existing military command systems and doctrine, so this approach is also less risky from a military procurement perspective.
@@ 4.3 Academic Policy and Action Recommendations
Lastly, we present several actions for academics and academic institutions to take to address the issues we describe in this work, beginning with an overview of recent actions researchers have taken on AWS regulation. Most recent academic initiatives have centered on open letters or petitions seeking to regulate or ban AWS (Future of Life Institute, 2016; Amoroso & Tamburrini, 2020). For example, in 2017 a Clearpath Robotics co-founder published an open letter calling for the UN to ban AWS, with 116 founders of robotics and AI companies from 26 countries signing on (Bogdon, 2017). Similarly, in 2018 AI researchers from 30 countries wrote an open letter protesting a new department at KAIST that merged academic and military research (Walsh, 2018). In 2018, Google employees’ protest led to the company canceling a Pentagon contract that could have led to more accurate drone strike technology (Crofts & van Rijswijk, 2020). Another Google employee protest has started recently, opposing their Project Nimbus (Perrigo, 2024). Furthermore, we note that global public opinion is favorable for AWS regulation currently. A 2020-2021 public opinion survey found that approximately 61 percent of adults across 28 countries oppose the use of fully autonomous weapons, with India being the only country surveyed with majority support for AWS (Ipsos, 2021).
@@ Acknowledgements
Last, some may argue that AWS will become essential for national security, and thus that it is unethical to discourage nations from developing AWS because they are needed to maintain military deterrence. More broadly, one could ask under what extremes may highly-autonomous AWS development and production be viewed as a necessary evil, and discouraging it therefore unethical? As a hypothetical example, an aggressive world power may refuse to limit fully-autonomous, human-free AWS development, and over years or decades their technology could progress to the point where it is impossible to counter using conventional forces, human-controlled AWS, or nuclear deterrence. In this particular situation we concede that fully-autonomous AWS development with a deterrence mentality, mirroring modern nuclear deterrence and tenuous geopolitical equilibriums, may pragmatically become a necessity. However, as outlined in this paper, while AWS battlefield deployments have started in preliminary ways, AWS are currently not the core of any nation’s military, and global public and institutional opinions are largely in favor of substantial regulation of some form. As such, this hypothetical is far from guaranteed to come to pass. However, we do not believe that the contributions of this paper are contingent on taking a position regarding the above hypothetical. Our work concerns the impacts of AWS on conflict risk and academic freedom, which will occur regardless of whether AWS are under strict human control or are highly autonomous, so long as they are critical to warfare and replace the human battlefield element. The goal of this work is to open discussion on the impacts of AWS among AI researchers and ground it in prior work, and we expect that the ethics of meaningful human control will be a major topic of future debate in the community, but our contribution does not rely on taking an ethical position on the matter.